What is the ratio of choosing a random red card out of a standard deck?
Since there are 26 red cards in all and there are 52 cards in a standard deck, the ratio would be, 26 to 52 or 1 to 2.
Is a standard deviation dependent on the mean?
To find the standard deviation, you must first compute the mean for the data set.
So the answer is yes. Just have a look at the 5 steps needed to compute a standard deviation and you will see why the answer is yes.
In reality, people most often use calculators or computers to do this.
However, it is good to understand what they are doing.
1. Compute the deviation by subtracting the mean from each value.
2. Square each individual deviation.
3. Add up the squared deviations.
4. Divide by one less than the sample size.
5. Take the square root
There is insufficient information for us to answer this question. Please edit the question to include more context or relevant information.
What is flood recurrence interval?
The Recurrence Interval is an estimate of the average time between past occurrences of random events. Typically, the period of record of these events must be at least 10 years to be statistically significant. Statistics are used to estimate the probability of the occurrence of some event. The following is from a USGS water science publication: "Statistical techniques, through a process called frequency analysis, are used to estimate the probability of the occurrence of a given precipitation event. The recurrence interval is based on the probability that the given event will be equalled or exceeded in any given year." The equation for figuring Recurrance Interval(RI) is: RI = (n+1)/m where n = number of years in the record, m = magnitude ranking For example; A flood event that had a peak stream discharge of 75,800 cfs ranked as the 3rd largest discharge over a 30 year record period. Therefore the RI is: RI = (30+1)/3 = 11yrs. ;meaning the likelihood of a flood event of that magnitude recurring within 11 years is fairly high. To estimate the exact probability of this event actually ocurring requires one to apply statistical analysis of probability; that is' P = 1 - (1 - 1/RI) (expression in parenthesis is raised to the X power where X is the number of years - 11 in this case)
How many jacks facing left are there in a deck of cards?
In a standard deck of 52 cards, the two red jacks face left.
What is the probability of getting a head or a tail when spinning 2 coins?
The "or" in the statement is addition; so even with one coin the probability of getting a head or a tail when spinning it is 1, and since the probability can never be greater than 1, the answer is 1.
Correct. I suspect the questioner means something different.
Probability of getting H & H = 1/2x1/2 = 1/4
Probability of getting H & T = 1/2 x 1/2 = 1/4
Probability of getting T & H = 1/2 x 1/2 = 1/4
Probability of getting T & T = 1/2 x 1/2 = 1/4
So probability of getting just 1 head = 1/4 +1/4 = 1/2
Probability of getting 2 heads = 1/4
So probability of getting 1 or 2 heads = 1/2 + 1/4 = 3/4
Probability of not getting a head = 1/4
Similarly for switching heads and tails.
What is the chances for having three sixes on a 3 dice?
The chances are (1/6) x (1/6) x (1/6)
The chances are (1/6) x (1/6) x (1/6)
The chances are (1/6) x (1/6) x (1/6)
The chances are (1/6) x (1/6) x (1/6)
What is the probability of choosing a queen from a standard deck playing cards?
4 in 52, which equals 1 in 13.
What is the beta distribution formula?
The probability density function for a beta distribution with parameters a and b (a, b > 0) is of the form:B(x; a, b) = c*x*(a-1)*(1-x)(b-1) where 0 <= x <= 1.
c is a constant such that the integral is 1 and it can be shown that c = G(a+b)/[G(a)*G(b)] where G is the Gamma function.
What is the probability of not rolling a 6 on a number cube?
The probability of not rolling it ever is 0.
For n rolls it is (5/6)n so
for 10 rolls it is 0.1615
for 20 rolls it is 2.608*10-2
for 100 rolls it is 1.207*10-8 and so on.
1 in 3
Probability is found by considering all possible outcomes but without actually testing them?
empirical
If you toss a coin into a liquid that in denser than the coin what would happen?
The coin would float.
What is the probability of getting heads when you toss a coin?
50% or 1/2.
There is 1 heads on a coin (numerator)
There are 2 sides on a coin (denominator)
WhaT Is thE probability of getting a sum of 13 with one roll of a pair of fair number cubes 1-6?
The probability is 0.
What is the probability of rolling a number that is divisible by 2 on a six sided die?
Evens
or 50/50
as either 2 or 4 or 6 (all divisible by 2)
or 1 or 3 or 5 (not divisible by 2)
What is the percentile rank if the raw score is 34?
It is not possible to convert a raw score into a percentile without knowing the distribution of scores and key parameters of the distribution. Since none of this information is provided, it is not possible to give a sensible answer.
How do you get probability peebles 2nd edition solution?
To obtain the solutions for "Probability Peebles, 2nd Edition," you can check the publisher's website for any supplementary materials or solutions manuals they may offer. Additionally, academic resources like university libraries or online educational platforms might provide access to these solutions. Alternatively, consider reaching out to classmates or study groups to discuss and work through the problems together.
What is the probability of picking a diamond or a 7 out of deck of 52 cards?
In a standard deck of 52 cards, there are 13 diamonds and 4 sevens. However, one of the sevens is also a diamond, so to avoid double counting, we calculate the total as 13 (diamonds) + 4 (sevens) - 1 (the seven of diamonds) = 16 favorable outcomes. Therefore, the probability of picking a diamond or a 7 is 16 out of 52, which simplifies to ( \frac{4}{13} ).